Gloria invested 60,000 pesos for 1 year, part at 9% and part at 4 1/2% simple interest. if the accounts earn 3,438.75 total interest, how much was invested at each rate?
Answer by mananth(16946)

Investment I 9.00% per annum ---x
Investment II 4.50% per annum ---y

x + y= 60000 ------------------------1
9.00% x + 4.50% y = $3,438.75
Multiply by 100
9 x + 4.5 y = $343,875.00

Multiply (1) by -9
we get

-9 x -9 y = -540000.00

Add this to (2)

0 x -4.5 y = -$196,125.00

divide by -4.5

y = $43,583.33
Balance $16,416.67
CHECK
$16,416.67 $1,477.50
$43,583.33 $1,961.25
Total $3,438.75